Fermi National Laboratory
 

Operation Summary for April 26, 2010 - April 28, 2010

 

Monday April 26

The day shift began with Operations monitoring the Tevatron (TeV) store 7768 (155.76E30), the Recycler stash (188.1E10), and the Antiproton Source stack (17.8E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, the NuMI kicker fluorinert temperature came into alarm - the fix was to open a secondary manual-cooling valve on the system to allow more flow through the heat exchanger. One antiproton transfer to the Recycler was lost when a Pbar Debuncher quadrupole (D:Q901) power supply faulted off; EE Support replaced a bad phase balance relay in the supply.

During the evening shift, Operations terminated store 7768 at 6:31 PM, put the TeV into shot setup, and established store 7769 at 7:39 PM with an average initial luminosity (AIL) of 391.34E30. Controls Group personnel replaced the clock decoder card for a Pbar Accumulator RF station (ARF4).

Tuesday April 27

The midnight shift began with Operations monitoring store 7769 (174.19E30) and the stash (128.4E10), with Pbar stacking (30.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Pbar stacked and transferred antiprotons to the Recycler. A Linac RF station (LRF4) tripped off; it reset.

During the day shift, Operations terminated store 7769 at 11:41 AM, put the TeV into shot setup, and established store 7770 at 1:09 PM with an AIL of 375.1E30. Water Group personnel made up LCW to the MI and TeV. FESS personnel worked on the MI-10 pond water pump, which interrupted beam for a few minutes a couple of times during the shift. Operators searched and secured Switchyard enclosures C, D, & E, which were followed by interlock safety system testing. Operations established slow spill extraction to the Meson beamline at 2:30 PM so that beamline physicists could tune the line. Operations received signoffs and permission to send beam to the MTest T977 MINERvA experiment. Operations established beam to MTest at 3:31 PM.

During the evening shift, Operations disabled the MTest beam at 6 PM; but reported some downtime due to a magnet interlock chassis trip for a Lambertson magnet power supply (MLAM1).

Wednesday April 28

The midnight shift began with Operations monitoring store 7770 (99.38E30) and the stash (293.1E10), with Pbar stacking (10.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac Klystron RF station (KRF3) tripped off; it reset. Operations established beam to MTest at 4 AM. Operations terminated store 7770 at 6:04 AM, put the TeV into shot setup, and established store 7771 at 7:24 AM with an AIL of 386.5E30.

The Plans for Wednesday

The plans for today are stack, stash, and store. The Run Coordinator will allow Linac work and a MI-31 access during the day shift, and a MI slow spill study this evening.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/28/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory